OpenStudy (anonymous):

you have 500,000 dollars explain each scenerio and decide which is the best one to use?? Option 1: 6% compounded interest quarterly for 5 years. Option 2: 8% compounded interest annually for 5 years. Option 3: 14.5% simple interest for 10 years.

OpenStudy (anonymous):

anyone

jimthompson5910 (jim_thompson5910):

For options 1 and 2, you'll use the formula A = P(1+r/n)^(n*t) For Option 1, you'll plug in P = 500000, r = 0.06, n = 4, and t = 5 For Option 2, you'll plug in P = 500000, r = 0.08, n = 1, and t = 5 ----------------------------------------------------------------- For option 3, you'll use the formula A = P(1+r*t) where in this case, P = 500000, r = 0.145, and t = 10

OpenStudy (anonymous):

ok could you put this in more simple terms? I understand princple* rate* time

jimthompson5910 (jim_thompson5910):

Simple Interest: Interest = princple* rate* time I = P*r*t with me so far?

OpenStudy (anonymous):

ok so im trying to do compound interset and yes so far lol thnks

jimthompson5910 (jim_thompson5910):

With a simple interest loan, you simply pay back the principle (the amount borrowed) plus some amount of interest (what we just found above)

jimthompson5910 (jim_thompson5910):

So if you are loaned P dollars, then you have to pay back P+I dollars

OpenStudy (anonymous):

what is n? i understand all the other terms

jimthompson5910 (jim_thompson5910):

n = compounding frequency

jimthompson5910 (jim_thompson5910):

n = 4 means you compound 4 times a year

jimthompson5910 (jim_thompson5910):

n = 1 means you compound once a year

OpenStudy (anonymous):

ok i kinda get it

jimthompson5910 (jim_thompson5910):

that's great

OpenStudy (anonymous):

so A = P(1+r/n)^(n*t) would be 500,000* 0.06/4 +4*4?

jimthompson5910 (jim_thompson5910):

A = P(1+r/n)^(n*t) A = 500000(1+0.06/4)^(4*5)

OpenStudy (anonymous):

ok so symbols confuse me, maybe you could right it out like 500000 times 1+ 0.06 and so on

jimthompson5910 (jim_thompson5910):

* means times example: 6*2 = 6 times 2

jimthompson5910 (jim_thompson5910):

so it's a handy shortcut